# On an approach to canonicalizing elliptic Feynman integrals

### v2

@@ 2.1 Basics for elliptic Feynman integrals
The incomplete elliptic integral of the first kind is defined as
where the differential is holomorphic. The second kind is defined as
@@
2.2 The construction of the pre-canonical basis
We can perform a variable change to transform the polynomial $P_{4}(z)$, which originates from the Baikov representation, to $P_{L}(x)$ as mentioned previously. The details of this transformation will be given later. After the transformation, the four branch points $\{c_{1},c_{2},c_{3},c_{4}\}$ defining the elliptic curve are transformed to $e_{1}=\infty$, $e_{2}=0$, $e_{3}=\lambda$ and $e_{4}=1$. We assume that the branch point $c_{i}$ for $4<i\leq n$ gets transformed to $e_{i}$, and $c_{n+1}=c_{\infty}$ is transformed to $e_{n+1}\equiv e_{\infty}$. We also want to stress that in $P_{L}(x)$, the point at the infinity or, say, $e_{1}$ is a branch point of the curve. However, $e_{\infty}$, which corresponds to the point of infinity in the original $P_{4}(z)$, is different from $e_{1}$ and no longer the point at infinity in the Legendre family. The transformed $u$ function is then given by
$u_{L}(x)=[P_{L}(x)]^{-1/2}\prod_{i=2}^{n+1}(x-e_{i})^{-\beta_{i}\varepsilon}\,,$ (13)
where
$\beta_{n+1}=-\sum_{i=1}^{n}\beta_{i}\,.$ (14)
We can then consider integrals of the form
[…]
where for later convenience we have defined
[…]
and absorb $P^{-1/2}_{4}(z)$ into the differential 1-form as the above.
@@ 2.2 The construction of the pre-canonical basis
$\phi_{1}$ $=\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=2\,{\rm d}F\left(\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)\,,$ (18a)
$\phi_{i-3}$ $=\frac{1}{x-e_{i}}\frac{\sqrt{P_{L}(e_{i})}}{\sqrt{P_{L}(x)}}{\rm d}x=-\frac{2\sqrt{P_{L}(e_{i})}}{e_{i}}\,{\rm d}\Pi\left(\frac{\lambda}{e_{i}},\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)\,,$ (18b)
$\phi_{n-1}$ $=\frac{x{\rm d}x}{\sqrt{P_{L}(x)}}=2\,{\rm d}F\left(\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)-2\,{\rm d}E\left(\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)\,,$ (18c)
@@
where $i=5,\cdots,n+1$. Note that the number of master integrals is $n-1$. The first and the last candidates involve differentials of the first and second kinds of standard elliptic integrals, while the rest involve the third kind. Plugging these $\phi_{j}$’s into Eq. (15), we get the corresponding master integrals $I_{j}$’s.
@@
We can now derive the differential equations of the master integrals with respect to $\lambda$ and $e_{i}$ for $i=5,\cdots,n+1$. We find that they are in a special linear form:
${\rm d}\vec{I}=\bm{A}\cdot\vec{I}=\left({\bm{A}^{(0)}}+\varepsilon{\bm{A}^{(1)}}\right)\cdot\vec{I}\,,$ (19)
@@
where the matrices $\bm{A}^{(0)}$ and ${\bm{A}^{(1)}}$ are independent of $\varepsilon$. To convert the above differential equations to the $\varepsilon$-factorized form, we only need to rotate $\bm{A}^{(0)}$ away by a linear transformation of the master integrals $\vec{I}$. Therefore, the explicit form of $\bm{A}^{(0)}$ is particularly important in this context. The coefficient matrix of $\mathrm{d}\lambda$ in $\bm{A}^{(0)}$ has the form
@@
$\begin{pmatrix}\frac{1}{2(1-\lambda)}&0&\cdots&0&-\frac{1}{2\lambda(1-\lambda)}\\ -\frac{e_{5}(e_{5}-1)}{2(1-\lambda)\sqrt{P_{L}(e_{5})}}&0&\cdots&0&\frac{e_{5}(e_{5}-1)}{2\lambda\left(1-\lambda\right)\sqrt{P_{L}(e_{5})}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ -\frac{e_{n+1}(e_{n+1}-1)}{2(1-\lambda)\sqrt{P_{L}(e_{n+1})}}&0&\cdots&0&\frac{e_{n+1}(e_{n+1}-1)}{2\lambda(1-\lambda)\sqrt{P_{L}(e_{n+1})}}\\ \frac{1}{2(1-\lambda)}&0&\cdots&0&-\frac{1}{2(1-\lambda)}\end{pmatrix},$ (20)
while the coefficient matrix of $\mathrm{d}e_{i}$ ($i=5,\cdots,n+1$) is
@@
$\begin{pmatrix}0&0&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ -\frac{e_{i}}{2\sqrt{P_{L}(e_{i})}}&0&\cdots&0&\frac{1}{2\sqrt{P_{L}(e_{i})}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&0\end{pmatrix}.$ (21)
The special linear form as Eq. (19) is an essential step towards deriving the $\varepsilon$-factorized form of the integral family, which will be elaborated in detail in Sec. 2.3 and App. A. In particular, $\bm{A}^{(0)}$ is closely related to the geometry behind the integral family and plays an important role in the later derivation. It is a good place here to put some remarks about it. For this purpose, we take $\varepsilon=0$ in this discussion. With a bit of abuse of notation, we are interested in the following equation:
${\rm d}\vec{I}^{(0)}={\bm{A}^{(0)}}\cdot\vec{I}^{(0)}\,.$ (22)
[…]
which equivalently defines the elliptic curve while
${\rm d}I_{i-3}^{(0)}=-\left[\frac{e_{i}(e_{i}-1)}{(1-\lambda)}\frac{{\rm d}\lambda}{2\sqrt{P_{L}(e_{i})}}+\frac{e_{i}{\rm d}e_{i}}{2\sqrt{P_{L}(e_{i})}}\right]I_{1}^{(0)}+\left[\frac{e_{i}(e_{i}-1)}{\lambda(1-\lambda)}\frac{{\rm d}\lambda}{2\sqrt{P_{L}(e_{i})}}+\frac{{\rm d}e_{i}}{2\sqrt{P_{L}(e_{i})}}\right]I_{n-1}^{(0)},$ (24)
with $15\leq i\leq n+$ 1, which is related to the marked points information in that curve. Denoting
$E=E\left(\sin^{-1}\sqrt{\frac{e_{i}}{\lambda}},\lambda\right),\quad F=F\left(\sin^{-1}\sqrt{\frac{e_{i}}{\lambda}},\lambda\right)$ (25)
${\rm d}I_{i-3}^{(0)}$ $=-\left[{\rm d}(F-E)-E\frac{{\rm d}\lambda}{2(1-\lambda)}\right]I_{1}^{(0)}+\left[{\rm d}F+\left(\frac{F}{2\lambda}-\frac{E}{2\lambda(1-\lambda)}\right){\rm d}\lambda\right]I_{n-1}^{(0)}$ (26)
$=-{\rm d}\left[(F-E)I_{1}^{(0)}-FI_{n-1}^{(0)}\right],$
where the second equality in the above follows with the help of the elliptic information (23). This property motivates us to refine the master integrals $I_{i-3}$ for $15\leq i\leq n+$ 1, which are directly related to the marked points as
$I_{i-3}\to I_{i-3}+(F-E)I_{1}-FI_{n-1}.$ (27)
2.3 The transformation to canonical form
Before we move to the explicit construction of a canonical basis, we would like to give short comments on the relationships between canonical bases and $\varepsilon$-factorized bases. $\varepsilon$-factorization is only one property for a canonical basis, and it cannot uniquely fix the basis, even up to a constant transformation. Such phenomena can stem in polylogarithmic cases as well, if we only ask for $\varepsilon$-factorization. To be more precise, the transformation to an $\varepsilon$-factorized basis can involve transcendental functions like logarithmic functions, while no such functions and only rational or algebraic functions are needed for the transformation to a canonical basis. This fact indicates that there should be some bases more “canonical” than others and it is natural to expect they would degenerate to canonical bases where the elliptic curves degenerate to Riemann spheres Frellesvig:2023iwr .
Our strategy to transform the pre-canonical basis to a canonical one is similar to that of Pogel:2022vat . We note that the method in Gorges:2023zgv leads to equivalent results. The essence of the strategy is to construct the canonical basis with ansatz inspired by the Picard-Fuchs operator for the integral of the first kind, $I_{1}$. In order to apply the method, we first introduce the periods of $\phi_{1}$, which serve as the solutions to the corresponding Picard-Fuchs equations. The periods are the pairings between the differential forms and the integration contours, i.e.,
$\varpi_{0}(\lambda)\equiv\frac{2}{\pi}\int_{0}^{\lambda}\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=\frac{4}{\pi}K(\lambda)\,,\quad\varpi_{1}(\lambda)\equiv\frac{2i}{\pi}\int_{\lambda}^{1}\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=\frac{4}{\pi}K(1-\lambda)\,,$ (28)
@@ 2.3 The transformation to canonical form
where $\varpi_{0}(\lambda)$ and $\varpi_{1}(\lambda)$ are holomorphic around degenerate points $\lambda=0$ and $\lambda=1$, respectively. The $1/\pi$ factors are introduced to make the periods of uniform transcendental weight-00 around degenerate points, while the $i$ factor in the second period is to make the periods real. [^6] Practically, we will pick different periods around different degenerate points, and thus we make the second period real to be more similar to the first one, which differs from most of the literature. The modular $\tau$ is defined as
$\tau\equiv\frac{i\varpi_{1}(\lambda)}{\varpi_{0}(\lambda)}=\frac{iK(1-\lambda)}{K(\lambda)}\in{\mathbb{H}}\,,$ (29)
where $\mathbb{H}$ is the upper-half complex plane.
The main idea of the ansatz is to normalize $I_{1}$ with its period, and then construct the integral of the second kind $\tilde{I}_{n-1}$ with the derivative basis of the normalized $\tilde{I}_{1}$ with respect to the modular variable $\tau$:
$\tilde{I}_{1}=\frac{1}{\varpi_{0}(\lambda)}I_{1}\,,\quad\tilde{I}_{n-1}=\frac{1}{\varepsilon}\frac{\varpi_{0}(\lambda)^{2}}{\pi W_{\lambda}}\partial_{\lambda}\tilde{I}_{1}\,,$ (30)
where $W_{\lambda}$ is the Wronskian of $\varpi_{0}$ and $\varpi_{1}$ with respect to $\lambda$
$W_{\lambda}=\varpi_{0}\partial_{\lambda}\varpi_{1}-\varpi_{1}\partial_{\lambda}\varpi_{0}=-\frac{4}{\pi\lambda(1-\lambda)}\,.$ (31)
With basis integrals in Eqs. (30) of the first and the second kind, we keep the rest of basis integrals (of the third kind) as pre-canonical basis integrals $I_{j}$ for $22\leq j\leq n-$ 2. We denote the transformation from pre-canonical basis $\vec{I}$ to $\tilde{\vec{I}}$ as ${\cal T_{\rm der}}$
$\tilde{\vec{I}}={\cal T}_{\rm der}\cdot\vec{I}\,,$ (32)
where the non-trivial elements are
$({\cal T}_{\rm der})_{1,1}$ $=\frac{1}{\varpi_{0}(\lambda)}\,,$ (33a)
$({\cal T}_{\rm der})_{n-1,n-1}$ $=\frac{1+2\beta_{1}\varepsilon}{8\varepsilon}\varpi_{0}(\lambda)\,,$ (33b)
$({\cal T}_{\rm der})_{n-1,j}$ $=-\frac{1}{4}\beta_{j+3}\frac{\sqrt{P_{L}(e_{j+3})}}{e_{i}-\lambda}\varpi_{0}(\lambda)\,,$ (33c)
@@
$({\cal T}_{\rm der})_{n-1,1}$ $=-\frac{1}{8\varepsilon}\left[\lambda\varpi_{0}(\lambda)-2\lambda(1-\lambda)\varpi_{0}^{\prime}(\lambda)+2\varepsilon\sum_{i=2}^{n+1}\beta_{i}(e_{i}-1)\varpi_{0}(\lambda)\right]\,,$ (33d)
where $22\leq j\leq n-$ 2 and other diagonal elements are equal to $1$. The corresponding connection matrix $\tilde{\bm{A}}$ is in a special linear form as Eq. (19), where the $\varepsilon^{0}$-part $\tilde{\bm{A}}^{(0)}$ is strictly lower triangular. The non-zero elements in the coefficient matrix of ${\rm d}e_{i}$ in $\tilde{\bm{A}}^{(0)}$ are given by [^7] The coefficient matrix of ${\rm d}\lambda$ is more complicated and is not so important for subsequent discussions. Therefore we do not show its entries here.
$(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}$ $=-\frac{(e_{i}-\lambda)\varpi_{0}(\lambda)+2\lambda(1-\lambda)\varpi_{0}^{\prime}(\lambda)}{2\sqrt{P_{L}(e_{i})}}\,,$ (34a)
$(\tilde{\bm{A}}^{(0)}_{e_{i}})_{n-1,1}$ $=\frac{\beta_{i}\lambda(1-\lambda)\varpi_{0}(\lambda)\varpi^{\prime}_{0}(\lambda)}{4(e_{i}-\lambda)}\,,$ (34b)
$(\tilde{\bm{A}}^{(0)}_{e_{i}})_{n-1,i-3}$ $=-\frac{\beta_{i}\lambda(1-\lambda)[\varpi_{0}(\lambda)-2(e_{i}-\lambda)\varpi^{\prime}_{0}(\lambda)]}{8(e_{i}-\lambda)\sqrt{P_{L}(e_{i})}}\,.$ (34c)
We first remove the $(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}$, which can be achieved by integrating it with respect to $e_{i}$. It turns out that the resulting function can be expressed with complete elliptic integral of the third kind. [^8] This fact is well-known Gorges:2023zgv ; Becchetti:2025rrz ; Becchetti:2025oyb . From the explicit form of $(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}$, it is also clear that the function we need can be expressed with products of complete and incomplete elliptic integrals of the first and second kinds, which indicates a non-trivial relation with complete elliptic integrals of the third kind. We conjecture a more formal derivation for this relation can be carried out with twisted Riemann bilinear relations Duhr:2024rxe . We will use complete integral of the third kind because of the compactness and its insensitivity with the choice of branch cut of $\sqrt{P_{L}(x)}$. Although we construct it from the connection matrix with respect to ${\rm d}e_{i}$, it also removes the $\varepsilon^{0}$-part with respect to $\lambda$. For simplicity, we will denote the resulting function as $\vartheta_{0}(e_{i},\lambda)$ hereafter, and it reads
which satisfies
$\frac{\partial\vartheta_{0}(e_{i},\lambda)}{\partial e_{i}}=-\frac{(e_{i}-\lambda)\varpi_{0}(\lambda)+2\lambda(1-\lambda)\varpi_{0}^{\prime}(\lambda)}{2\sqrt{P_{L}(e_{i})}}=(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}\,.$ (36)
We add a subscript “00” for it because it is holomorphic around $\lambda=0$ while logarithmic around $\lambda=1$, similar to $\varpi_{0}(\lambda)$. We can also define its counterpart $\vartheta_{1}(e_{i},\lambda)$ which shares similar asymptotic behaviors with $\varpi_{1}(\lambda)$
$\vartheta_{1}(e_{i},\lambda)\equiv-\frac{4\sqrt{P_{L}(e_{i})}}{\pi e_{i}}\left[K(1-\lambda)+\frac{\lambda}{e_{i}-\lambda}\Pi\left(\frac{e_{i}(1-\lambda)}{e_{i}-\lambda},1-\lambda\right)\right]\,,$ (37)
and it satisfies the same differential equation Eq. (36) as $\vartheta_{0}(e_{i},\lambda)$, with the subscript “00” substituted by “$1$”. This function is useful if one considers the region around $\lambda=1$.
We also need to remove $\varepsilon^{0}$-part for the last integral $\tilde{I}_{n-1}$, which is the most complicated one. Thanks to the method from Duhr:2024uid , we can make an ansatz that the final canonical basis has a constant intersection matrix, and then the ansatz transformation can be found directly. We can then give the transformation ${\cal T}$ from the pre-canonical basis $\tilde{\vec{I}}$, where the non-trivial elements are
${\cal T}_{j,1}$ $=-\vartheta_{0}(e_{j+3},\lambda)\,,$ (38a)
${\cal T}_{n-1,j}$ $=\frac{1}{4}\beta_{j+3}\left[\vartheta_{0}(e_{j+3},\lambda)+\frac{e_{j+3}(e_{j+3}-1)\varpi_{0}(\lambda)}{\sqrt{P_{L}(e_{j+3})}}\right]\,,$ (38b)
@@
${\cal T}_{n-1,1}$ $=-\frac{1}{8}\left[\sum_{i=5}^{n+1}\beta_{i}\vartheta_{0}(e_{i},\lambda)^{2}-\sum_{i=2}^{n+1}\beta_{i}(e_{i}+\lambda-1)\varpi_{0}(\lambda)^{2}\right]\,,$ (38c)
where diagonal elements are equal to $1$ and $22\leq j\leq n-$ 2. With ${\cal T}_{\rm der}$ from (33) and ${\cal T}$ from (38), the transformation from pre-canonical basis $\vec{I}$ to canonical basis ${\cal T}_{\rm can}$ can be obtained directly
${\cal T}_{\rm can}={\cal T}\cdot{\cal T}_{\rm der}\,.$ (39)
Now we take a look at the canonical connection matrix with respect to ${\rm d}\lambda$ near the degenerate point $\lambda=0$:

### v3

@@ 2.2 The construction of the pre-canonical basis with the Legendre normal form
The incomplete elliptic integral of the second kind is defined as
@@
$\phi_{1}^{\text{pre}}$ $=\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=2\,{\rm d}F\left(\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)\,$, (21a)
$\phi_{i-3}^{\text{pre}}$ $=\frac{1}{x-e_{i}}\frac{\sqrt{P_{L}(e_{i})}}{\sqrt{P_{L}(x)}}{\rm d}x=-\frac{2\sqrt{P_{L}(e_{i})}}{e_{i}}\,{\rm d}\Pi\left(\frac{\lambda}{e_{i}},\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)\,$, (21b)
$\phi_{n-1}^{\text{pre}}$ $=\frac{x{\rm d}x}{\sqrt{P_{L}(x)}}=2\,{\rm d}F\left(\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)-2\,{\rm d}E\left(\sin^{-1}\sqrt{\frac{x}{\lambda}},\lambda\right)\,$, (21c)
@@
where $i=5,\cdots,n+1$. Note that the number of master integrals is n1n-1. The first and the last candidates involve differentials of the first and second kinds of standard elliptic integrals, while the rest involve the third kind. Plugging these $\phi_{j}^{\text{pre}}$ into Eq. (15), we get the corresponding master integrals $I_{j}^{\text{pre}}$. Note we do not introduce any elliptic integrals into the integrands at this step.
@@
We can now derive the differential equations of the pre-canonical master integrals with respect to $\lambda$ and $e_{i}$ for $i=5,\cdots,n+1$. We find that they are in a linear form:
${\rm d}\vec{I}^{\text{pre}}=\bm{A}\,\vec{I}^{\text{pre}}=\left({\bm{A}^{(0)}}+\varepsilon{\bm{A}^{(1)}}\right)\vec{I}^{\text{pre}}\,$, ($2$)
@@
where the matrices $\bm{A}^{(0)}$ and $\bm{A}^{(1)}$} are independent of $\varepsilon$. To convert the above differential equations to the $\varepsilon$-factorized form, we only need to rotate the matrix $\bm{A}$ by a linear transformation of the master integrals $\vec{I}^{\text{pre}}$, such that the $\varepsilon^{0}$-part is absent. Before doing that, we may have a look at the explicit form of $\bm{A}^{(0)}$. We write $\bm{A}^{(0)}=\bm{A}^{(0)}_{\lambda}\mathrm{d}\lambda+\sum_{i}\bm{A}^{(0)}_{e_{i}}\mathrm{d}e_{i}$, where
@@
$\bm{A}^{(0)}_{\lambda}=\begin{pmatrix}\frac{1}{2(1-\lambda)}&0&\cdots&0&-\frac{1}{2\lambda(1-\lambda)}\\ -\frac{e_{5}(e_{5}-1)}{2(1-\lambda)\sqrt{P_{L}(e_{5})}}&0&\cdots&0&\frac{e_{5}(e_{5}-1)}{2\lambda\left(1-\lambda\right)\sqrt{P_{L}(e_{5})}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ -\frac{e_{n+1}(e_{n+1}-1)}{2(1-\lambda)\sqrt{P_{L}(e_{n+1})}}&0&\cdots&0&\frac{e_{n+1}(e_{n+1}-1)}{2\lambda(1-\lambda)\sqrt{P_{L}(e_{n+1})}}\\ \frac{1}{2(1-\lambda)}&0&\cdots&0&-\frac{1}{2(1-\lambda)}\end{pmatrix}\,$, (23)
and
@@
$\bm{A}^{(0)}_{e_{i}}=\begin{pmatrix}0&0&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ -\frac{e_{i}}{2\sqrt{P_{L}(e_{i})}}&0&\cdots&0&\frac{1}{2\sqrt{P_{L}(e_{i})}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&0\end{pmatrix}\,,\quad(i=5,\cdots,n+1)\,.$ (24)
Notice $\bm{A}^{(0)}$ is special where only the first and the last columns are non-vanishing. This is a direct consequence of the specific choice of the pre-canonical basis in Eq. (21). This fact will be important when we consider the sub-sector dependence of the canonical basis.
2.3 The transformation to the canonical basis
We now want to turn the pre-canonical integrands $\vec{\phi}^{\text{pre}}$ in (21) into a canonical one by acting onto it a transformation matrix ${\cal T}_{\rm can}$. We will perform the transformation in two steps. In other words, we split the matrix ${\cal T}_{\rm can}$ into two parts:
${\cal T}_{\rm can}={\cal T}\,{\cal T}^{\prime}\,.$ (25)
The matrix $\cal T$$^{\prime}$ changes only the first and the last integrals. It is related to the periods of the elliptic curve, defined as the pairings between the differential forms and the integration contours:
$\varpi_{0}(\lambda)\equiv\frac{2}{\pi}\int\limits_{0}^{\lambda}\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=\frac{4}{\pi}K(\lambda)\,,\quad\varpi_{1}(\lambda)\equiv\frac{2i}{\pi}\int\limits_{\lambda}^{1}\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=\frac{4}{\pi}K(1-\lambda)\,$, (26)
@@ 2.3 The transformation to the canonical basis
where $\varpi_{0}(\lambda)$ and $\varpi_{1}(\lambda)$ are holomorphic around degenerate points $\lambda=0$ and $\lambda=1$, respectively. The $1/\pi$ factors are introduced to make the periods of uniform transcendental weight-0 around degenerate points, while the $i$ factor in the second period is to make the periods real. [^5] The convention here is slightly different from many literatures.
The periods are solutions to the Picard-Fuchs equation satisfied by the first pre-canonical integral $I_{1}^{\text{pre}}$. It is therefore natural to normalize $I_{1}^{\text{pre}}$ by one of the periods. Here we choose $\varpi_{1}(\lambda)$, which is holomorphic as $\lambda\to 1$. As for $I_{n-1}^{\text{pre}}$, we will rotate it to the derivative of the normalized $I_{1}^{\text{pre}}$ with respect to the modular variable $\tau$ defined as (in the region around $\lambda\to 1$):
$\tau\equiv\frac{i\varpi_{0}(\lambda)}{\varpi_{1}(\lambda)}=\frac{iK(\lambda)}{K(1-\lambda)}\in{\mathbb{H}}\,$, (27)
where $\mathbb{H}$ is the upper-half complex plane. The above strategy is similar to that of Pogel:2022vat , which is also equivalent to Gorges:2023zgv . Nevertheless, we work with a generic elliptic family in the Legendre normal form rather than a particular Feynman integral family, and our goal is to derive the generic result (10) of the canonical basis after a Möbius transformation.
In summary, the non-vanishing elements of the matrix $\cal T$$^{\prime}$ are given by
$({\cal T}^{\prime})_{1,1}$ $=\frac{1}{\varpi_{1}(\lambda)}\,$, (28a)
$({\cal T}^{\prime})_{n-1,n-1}$ $=-\frac{1+2\beta_{1}\varepsilon}{8\varepsilon}\varpi_{1}(\lambda)\,$, (28b)
$({\cal T}^{\prime})_{n-1,j}$ $=\frac{1}{4}\beta_{j+3}\frac{\sqrt{P_{L}(e_{j+3})}}{e_{i}-\lambda}\varpi_{1}(\lambda)\,,\quad({\cal T}^{\prime})_{j,j}=1\,,\quad(2\leq j\leq n-2)\,$, (28c)
@@
$({\cal T}^{\prime})_{n-1,1}$ $=\frac{1}{8\varepsilon}\left[\lambda\varpi_{1}(\lambda)-2\lambda(1-\lambda)\varpi_{1}^{\prime}(\lambda)+2\varepsilon\sum_{i=2}^{n+1}\beta_{i}(e_{i}-1)\varpi_{1}(\lambda)\right]\,.$ (28d)
The matrix transforms $I_{n-1}^{\text{pre}}$ into
@@
$\left({\cal T}^{\prime}\vec{I}^{\text{pre}}\right)_{n-1}=\frac{1}{\varepsilon}\frac{\varpi_{1}(\lambda)^{2}}{\pi W_{\lambda}}\partial_{\lambda}\bigg[\frac{1}{\varpi_{1}(\lambda)}I_{1}^{\text{pre}}\bigg]\,$, (29)
where $W_{\lambda}$ is the Wronskian of $\varpi_{1}$ and $\varpi_{0}$ with respect to $\lambda$
$W_{\lambda}=\varpi_{1}\partial_{\lambda}\varpi_{0}-\varpi_{0}\partial_{\lambda}\varpi_{1}=\frac{4}{\pi\lambda(1-\lambda)}\,.$ (30)
After imposing $\cal T$$^{\prime}$, the connection matrix in the differential equations of ${\cal T}^{\prime}\vec{I}^{\text{pre}}$ is still in a linear form as Eq. (22). We denote the $\varepsilon^{0}$-part of the connection matrix as $\tilde{\bm{A}}^{(0)}=\tilde{\bm{A}}^{(0)}_{\lambda}\mathrm{d}\lambda+\sum_{i}\tilde{\bm{A}}^{(0)}_{e_{i}}\mathrm{d}e_{i}$, which is strictly lower triangular. The non-zero elements in $\tilde{\bm{A}}^{(0)}_{e_{i}}$ are given by [^6] The structure of the matrix $\tilde{\bm{A}}^{(0)}_{\lambda}$ is similar, albeit with more complicated entries. We do not show its entries here explicitly.
$(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}$ $=-\frac{(e_{i}-\lambda)\varpi_{1}(\lambda)+2\lambda(1-\lambda)\varpi_{1}^{\prime}(\lambda)}{2\sqrt{P_{L}(e_{i})}}\,$, (31a)
$(\tilde{\bm{A}}^{(0)}_{e_{i}})_{n-1,1}$ $=-\frac{\beta_{i}\lambda(1-\lambda)\varpi_{1}(\lambda)\varpi_{1}^{\prime}(\lambda)}{4(e_{i}-\lambda)}\,$, (31b)
$(\tilde{\bm{A}}^{(0)}_{e_{i}})_{n-1,i-3}$ $=\frac{\beta_{i}\lambda(1-\lambda)[\varpi_{1}(\lambda)-2(e_{i}-\lambda)\varpi_{1}^{\prime}(\lambda)]}{8(e_{i}-\lambda)\sqrt{P_{L}(e_{i})}}\,.$ (31c)
We now need to rotate these entries away with the matrix $\mathcal{T}$ in Eq. (25).
We first construct the matrix element that removes $(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}$. It can be obtained by solving the differential equation
$\frac{\partial\vartheta_{1}(e_{i},\lambda)}{\partial e_{i}}=(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}=-\frac{(e_{i}-\lambda)\varpi_{1}(\lambda)+2\lambda(1-\lambda)\varpi_{1}^{\prime}(\lambda)}{2\sqrt{P_{L}(e_{i})}}\,.$ (32)
We denote the function as $\vartheta_{1}$ to indicate that it is holomorphic around $\lambda=1$, similar to $\varpi_{1}(\lambda)$. It turns out that the solution can be expressed with the complete elliptic integral of the third kind Gorges:2023zgv ; Becchetti:2025rrz ; Becchetti:2025oyb : [^7] By directly integrating $(\tilde{\bm{A}}^{(0)}_{e_{i}})_{i-3,1}$, it is also clear that $\vartheta_{1}(e_{i},\lambda)$ can be expressed in terms of the complete and incomplete elliptic integrals of the first and second kinds. This indicates a non-trivial relation among elliptic integrals, which might be provable with the twisted Riemann bilinear relations Duhr:2024rxe .
$\vartheta_{1}(e_{i},\lambda)=-\frac{4\sqrt{P_{L}(e_{i})}}{\pi e_{i}}\left[K(1-\lambda)+\frac{\lambda}{e_{i}-\lambda}\Pi\left(\frac{e_{i}(1-\lambda)}{e_{i}-\lambda},1-\lambda\right)\right]\,$, ($3$)
Although we have constructed $\vartheta_{1}$ from $\tilde{\bm{A}}^{(0)}_{e_{i}}$, it also removes the relevant entries in $\tilde{\bm{A}}^{(0)}_{\lambda}$, without the need to add further $\lambda$-dependent terms.
We still need to remove the entries in the last row of $\tilde{\bm{A}}^{(0)}_{e_{i}}$. We adopt the method of Ref. Duhr:2024uid , by looking for a transformation that leads to a constant intersection matrix for the canonical basis. Such a transformation can be easily found with an ansatz in terms of $\varpi_{1}$ and $\vartheta_{1}$ functions. As a result, we find that all diagonal elements of $\mathcal{T}$ are equal to 1, and the other non-zero elements are given by
${\cal T}_{j,1}$ $=-\vartheta_{1}(e_{j+3},\lambda)\,$, (34a)
${\cal T}_{n-1,j}$ $=-\frac{1}{4}\beta_{j+3}\left[\vartheta_{1}(e_{j+3},\lambda)+\frac{e_{j+3}(e_{j+3}-1)\varpi_{1}(\lambda)}{\sqrt{P_{L}(e_{j+3})}}\right]\,$, (34b)
${\cal T}_{n-1,1}$ $=\frac{1}{8}\left[\sum_{i=5}^{n+1}\beta_{i}\vartheta_{1}(e_{i},\lambda)^{2}-\sum_{i=2}^{n+1}\beta_{i}(e_{i}+\lambda-1)\varpi_{1}(\lambda)^{2}\right]\,$, (34c)
where $2\leq j\leq n-2$. We can now combine $\cal T$$^{\prime}$ from (28) and $\cal T$} from (34) into ${\cal T}_{\rm can}$ as in Eq. (25). Acting the matrix ${\cal T}_{\rm can}$ onto (21), we readily obtain a canonical basis for the integral family in the Legendre normal form.
It is interesting to have a look at the differential equations satisfied by our canonical basis. The connection matrix of the canonical basis has the $\varepsilon$-factorized form
$\bm{A}^{\rm can}={\bm{A}}_{\lambda}^{\rm can}{\rm d}\lambda+\sum_{i}{\bm{A}}_{e_{i}}^{\rm can}{\rm d}e_{i}=\varepsilon\bm{A}^{\text{can},(1)}\,.$ (35)
We find that the entries in $\bm{A}^{\rm can}$ only involve the transcendental functions $\varpi_{1}(\lambda)$ and $\vartheta_{1}(e_{i},\lambda)$, while $\varpi^{\prime}_{1}(\lambda)$ does not appear. The power of $\varpi_{1}(\lambda)$ in the denominator is at most $2$, and $\vartheta_{1}(e_{i},\lambda)$ only appears in the numerator. Furthermore, the matrix at the degenerate point $\lambda=1$ is UT and has only a simple pole, e.g.,